### v1

# Ultra precise determination of Cs($nS_{1/2}$) and Cs($nD_{J}$) quantum defects for sensing and computing: Evaluation of core contributions

Abstract
We make absolute frequency measurements of Cs Rydberg transitions, $|6S_{1/2},F=3\rangle\rightarrow|nS_{1/2}$ $(n=23\rm{-}90)\rangle$ and $|nD_{3/2,5/2}$ $(n=21\rm{-}90)\rangle$, with an accuracy of less than $72\,\rm kHz$. The quantum defect parameters for the measured Rydberg series are the most precise obtained to date. The quantum defect series is terminated at $\delta_{4}$, showing that prior fits requiring higher order quantum defects reflect uncertainties in the observations. The precision of the measured quantum defects allow for the calculation of Rydberg electric-dipole transitions and fine-structure intervals extrapolated from high principal quantum numbers, to rival that of sophisticated many-body relativistic calculations carried out at low Rydberg principal quantum numbers. We quantitatively predict the contributions to the quantum defect parameters from core polarization and core penetration of Cs inner shell electrons. A new value for the ionization energy, consistent across the $nS_{1/2}$ and $nD_{3/2,5/2}$ Rydberg series, is reported at $31406.46775148(14)~{}\rm{cm}^{-1}$.
Neutral-atom-based quantum simulators and processors are the leading architectures for high fidelity quantum operations while atoms are being used extensively as sensors because of their stability and uniformity. Rydberg atoms are particularly suitable for quantum simulations and metrology, because their properties can be tuned through state selection and the use of electromagnetic fields. Precise knowledge of atomic structure enables engineering of atomic properties. Similarly for basic research, precise measurements of atomic energy levels provide stringent tests of atomic structure calculations, including quantum defect theories, and are critical for the determination of a number of fundamental constants in physics [1, 2, 3]. Cesium atoms have been used for searches of the electric-dipole moment of the electron [4, 5, 6, 7], amongst other fundamental quantities. Rapid developments in quantum measurement technologies, such as optical clocks [8, 9, 10, 11, 12], neutral-atom-based quantum computing [13, 14, 15, 16, 17] and Rydberg-atom-based electric field sensors [18, 19, 20] are adding impetus to accurately determine energy levels as well as derived quantities such as electric field polarizabilities and reduced electric-dipole matrix elements. The aim of this work is to improve the accuracy of Cs atomic energy level measurements for advanced applications in quantum sensing.
We present absolute frequency measurements for the Cs transitions $|6S_{1/2},F=3\rangle\rightarrow|nS_{1/2}$ $(n=23\rm{-}90)\rangle$ and $|nD_{3/2,5/2}$ $(n=21\rm{-}90)\rangle$ with an accuracy of $<~{}72\,\rm kHz$. By fitting the absolute-frequency measurements to the modified Ritz formula, we determine the most accurate quantum defect parameters of the $nS_{1/2}$, $nD_{3/2}$, and $nD_{5/2}$ series to date, and obtain a new ionization energy, $31406.46775148(14)~{}\rm{cm}^{-1}$. Because of the accuracy of the underlying term energies, the highest order fitted quantum defect is $\delta_{4}$, illustrating that the higher terms, often found in the literature, reflect uncertainties in the measurements. We extract the fine-structure intervals for the $nD_{J}$ series. The precision of our calculations of electric-dipole transition matrix elements is comparable with that of sophisticated many-body relativistic calculations [21] at low $n$, where measurements were not carried-out. We calculate the contributions to the quantum defect parameters from core polarization and core penetration. Both effects are found to be in agreement with high level atomic structure calculations.
The experimental cycle begins by collecting Cs atoms in a magneto-optical trap (MOT). The MOT atoms are cooled to $<10~{}\mu\rm K$ using polarization gradient cooling [22, 23]. After the cooling stage, the atoms are optically pumped to the $|6S_{1/2},F=3\rangle$ ground state. Subsequently, the trapping magnetic field and lasers are switched off to create a field-free environment for the atoms. During the entire interrogation period, the residual magnetic field is reduced to $<2~{}\rm mG$ and the stray electric field is reduced to $<1$ mV/cm, using external field compensation [24].
A two-step laser excitation at $\sim 852~{}\rm nm$ and $\sim 509~{}\rm nm$ creates cold Cs Rydberg atoms. The lasers are both locked to a frequency comb that is stabilized to a GPS steered rubidium clock, resulting in laser linewidths of less than 200 Hz and a fractional frequency stability of $1.2\times 10^{-12}$ for a $10~{}\rm s$ averaging time. The $\sim 852~{}\rm nm$ laser is red-detuned 90 MHz from the $|6P_{3/2}$, $F^{\prime}=2\rangle$ excited state. The $\sim 509~{}\rm nm$ laser is 90 MHz blue-detuned from the resonance between the $|6P_{3/2}$, $F^{\prime}=2\rangle$ state and the target $nD_{J}$ or $nS_{1/2}$ Rydberg state. The two lasers are crossed in a perpendicular configuration, both with parallel linear, vertical polarizations. The Gaussian beam waists for the first- and second-step lasers are $w_{0}=3~{}\rm{mm}$ and $w_{0}=400~{}\mu\rm{m}$, respectively. The lasers are pulsed to excite the Rydberg atoms. The duration of the excitation laser pulses is $30~{}\mu$s. The pulses overlap in time.
Following the laser excitation pulses, a 2 $\mu$s high voltage square pulse is applied to field ionization plates to ionize the Rydberg atoms and project the ions onto a multi-channel plate (MCP) detector. The average number of detected ions generated per laser shot is minimized to $<2$ to reduce Coulomb repulsion. The ion pulses are recorded on a multichannel analyzer.
To obtain the Rydberg state spectrum, the frequency of the $\sim 852~{}\rm nm$ laser is fixed and the frequency of the $\sim 509~{}\rm nm$ laser is scanned. The laser excitation and ion detection sequence is repeated 40 times for each cold atomic cloud. Twenty clouds are used for each $\sim 509~{}\rm nm$ laser frequency, resulting in 800 excitation-detection cycles per laser frequency step. The spectroscopy measurements are performed for the $nD_{5/2}$ ($n=21$–90), $nD_{3/2}$ ($n=21$–90), and $nS_{1/2}$ ($n=23$–90) series.
For the $nD_{3/2}$ and $nD_{5/2}$ series, the measured spectrum is fit to a hyperfine-structure-weighted multiline Gaussian function [24]. For the $nS_{1/2}$ series spectrum, a single Gaussian function is used since there is only one allowed hyperfine transition, $|6P_{3/2},F=2\rangle\rightarrow|nS_{1/2},F^{\prime}=3\rangle$, and the dominant broadening mechanism is Doppler broadening, at the $200~{}\rm{kHz}$ level. The spectral line frequency is the absolute transition frequency between the $|6S_{1/2},F=3\rangle$ hyperfine state and the center of gravity of the Rydberg energy level.
The transition frequencies are corrected to account for the AC Stark shifts caused by the lasers. The AC Stark shift induced by the $\sim~{}509~{}\rm nm$ laser is no greater than $500~{}\rm{Hz}$ for the $21D_{J}$ state. The $\sim~{}852~{}\rm nm$ laser induces a measured ($36.0\pm 4.8)~{}\rm{kHz}$ blue shift of the ground state, $|6S_{1/2},F=3\rangle$ [24]. The 36.0 kHz shift is subtracted from the transition frequency and the uncertainty, ±$\pm 4.8~{}\rm kHz$, is taken as a contribution to the uncertainty of the resonance frequency.
[…]
The Zeeman shift of the ground state is minimized using microwave spectroscopy [24, 44]. The Zeeman shift was measured to be $<14.3~{}\rm kHz$. The sensitivity to the magnetic field is similar across all Rydberg states because they have the same Landé $g$-factor. The Zeeman shift is treated as another source of uncertainty.
The shot-to-shot variance dominates the statistical uncertainty. Systematic shifts, most probably associated with residual Stark shifts, can be observed in the plot of the residuals, Fig. 1. The sum of the statistical uncertainty and the systematic uncertainty is no greater than 60 kHz [24]. The statistical standard deviation of the residuals of the modified Ritz formula fit is $<72~{}\rm kHz$ for all series.
Figure 1: Absolute frequency measurements for transitions from $|6S_{1/2},F=3\rangle$ to the center of gravity of the $nS_{1/2}$, $nD_{3/2}$, $nD_{5/2}$ Rydberg states. Data are presented with a step of $\Delta n=3$ to avoid crowding. The dashed lines correspond to the modified Ritz formula. The fitting residuals are shown in the inset. The error bars represent the total uncertainty. The statistical variance of the fitting residuals of the $nD_{5/2}$, $nD_{3/2}$, and $nS_{1/2}$ series are $71.2~{}\rm kHz$, $59.1~{}\rm kHz$, and $46.9~{}\rm kHz$, respectively.
Table 1: Quantum defects for the $nD_{5/2}$, $nD_{3/2}$, and $nS_{1/2}$ series obtained by fitting the measured Rydberg energy levels to Eq. (1). The uncertainties are the standard deviations of the fit. For comparison, the quantum defects used in the ARC python calculation package [40] are also listed.
$nD_{5/2}$ [this work] $nD_{5/2}$ [45] $nD_{3/2}$ [this work] $nD_{3/2}$ [46] $nS_{1/2}$ [this work] $nS_{1/2}$ [45]
($n=21$–90) ($n=5$–36) ($n=21$–90) ($n=5,6,8$–32) ($n=23$–90) ($n=6$–30)
IE $(\rm{cm}^{-1})$ 31406.46775168(26) 31406.46767(3) 31406.46775148(28) 31406.4710(7) 31406.46775136(21) 31406.46769(2)
$\delta_{0}$ 2.46631529(14) 2.46631524(63) 2.47545840(13) 2.4754562 4.04935994(17) 4.04935665(38)
$\delta_{2}$ 0.013431(65) 0.013577 0.008339(62) 0.009320 0.238017(18) 0.2377037
$\delta_{4}$-0.3613(61) --0.37457 --0.3769(95) --0.43498 0.1747(33) 0.255401
$\delta_{6}$ 0.0 --2.1867 0.0 --0.76358 0.0 0.00378
$\delta_{8}$ 0.0 --1.5532 0.0 --18.0061 0.0 0.25486
$\delta_{10}$ 0.0 --56.6739 0.0 0.0 0.0 0.0
Table 2: Numerically computed core polarization contribution, $\delta_{\rm pol}^{(l,J)}$, and core penetration contribution, $\delta_{\rm pen}^{(l,J)}$, to the leading quantum defect term of the $nD_{5/2}$, $nD_{3/2}$, and $nS_{1/2}$ series.
series $\delta_{0}^{(l,J)}$ $\delta_{\rm pol}^{(l,J)}$ $\delta_{\rm pen}^{(l,J)}$
$nD_{5/2}$ 2.46631529(14) 1.0482(4) 1.4197(6)
$nD_{3/2}$ 2.47545840(13) 1.0625(7) 1.4094(5)
$nS_{1/2}$ 4.04935994(17) 0.1320(5) 3.9111(4)
The observed term energies of a single Rydberg series can be described by the modified Ritz formula:
$E_{n,l,J}=E_{I}-\frac{R}{[n-\delta^{(l,J)}(n)]^{2}},$ (1)
where $E_{I}$ is the ionization energy and $R$ is the reduced mass Rydberg constant, equal to $R_{\infty}/(1+m_{\rm e}/m_{\rm core})$. Using the current mass of the Cs atom [47], we obtained $R_{{}^{133}\rm Cs}=$ 109736.8627304 cm-1. The $(n,l,J)$-dependent quantum defects are
$\delta^{(l,J)}(n)=\delta^{(l,J)}_{0}+\sum_{k=1}^{\infty}\frac{\delta^{(l,J)}_{{2k}}}{[n-\delta^{(l,J)}_{0}]^{2k}}.$ (2)
We constrain the quantum defects to a three-term expansion, $\delta^{(l,J)}(n)=\delta^{(l,J)}_{0}+\frac{\delta^{(l,J)}_{{2}}}{[n-\delta^{(l,J)}_{0}]^{2}}+\frac{\delta^{(l,J)}_{{4}}}{[n-\delta^{(l,J)}_{0}]^{4}}$, and fit the measured energy levels to Eq. (1), which minimizes the correlations between the ionization energy and the quantum defects present in the fit. The Rydberg state energy is obtained by subtracting the frequency splitting, $5.170855370625~{}\rm GHz$, between the center of gravity of the $6S_{1/2}$ state and the $|6S_{1/2},F=3\rangle$ hyperfine state [48] from the measured absolute transition frequencies listed in Table S4 [24], i.e., the energies are referenced to the ground state center of gravity to perform the fitting.
The ionization energies are reported in Table 1. The extracted ionization energies for the $nD_{5/2}$, $nD_{3/2}$, and $nS_{1/2}$ states agree within their uncertainties. The uncertainty-weighted average ionization energy is $31406.46775148(14)~{}\rm{cm}^{-1}$. The result agrees with the ionization energy reported in Ref. [45], but is $600~{}\rm kHz$ higher than the most recent value presented in Ref. [49], $31406.4677325(14)~{}\rm{cm}^{-1}$. In Ref. [49], Cs samples were prepared in an optical dipole trap that cannot be completely shut-off, leading to a larger AC Stark effect. The ionization energy reported in Ref. [49] was extracted using one Rydberg series, $|nP_{J}$ $(n=27\rm{-}74)\rangle$ where the $n$ step size was 3. In this work, the Cs samples are excited in a field-free environment and the ionization energy is extracted from all of the levels in the three Rydberg series $|nS_{1/2}$ $(n=23\rm{-}90)\rangle$ and $|nD_{J}$ $(n=21\rm{-}90)\rangle$. Moreover, the uncertainty of our ionization energy determined by the fit to the modified Ritz formula is further reduced by one order of magnitude.
The non-linear least squares fit to measured energies are shown in Fig. 1 with residues in the inset. The systematic downward trend for the higher principal quantum numbers of the $nD_{J}$ series is likely caused by DC Stark shifts. The quantum defects for all three series are presented in Table 1. Our results show that the $\delta_{2k}^{(l,J)}$ parameters for $k>2$ for all three series are insignificant within the measurement uncertainties. Several studies have reported these terms for the same states as non-zero at lower levels of experimental precision [45, 46]. Since the precision of our measurements is two orders of magnitude better than that of the prior work, it is likely that these terms have no physical significance and were fitting parameters that accommodated experimental uncertainties.
Figure 2: Experimental quantum defects, $\delta^{(l,J)}$, versus $n$ (a) and $1/[n-\delta^{(l,J)}]^{2}$ (b), for all measured term series in Cs. Solid markers are the quantum defects measured in this work, while open markers are extrapolated from the experimental measurements. The light color markers represent the values calculated based on previous publications, $nS_{1/2}$ [45], $nD_{3/2}$ [46], and $nD_{5/2}$ [45]. The error bars are acquired by propagating the fitting uncertainty of the quantum defects.
The quantum defect parameters obtained from the measurements in Fig. 2(a) agree well within the uncertainties, even for lower-lying states ($218<n<$ 21), where no measurements were taken. The plots show further evidence that the quantum defect expansion greater than third order is not required, since our predictions for the $nS_{1/2}$, $nD_{3/2}$, and $nD_{5/2}$ energy levels match the available experimental data at low $n$. Notably, the trend for both $nD_{J}$ series is non-monotonic while that of the $nS_{1/2}$ series always decreases as a function of $n$. The non-monotonic behavior of the $nD_{J}$ series had been reported in Refs. [50, 46, 45]. The behavior can be viewed more clearly by plotting the quantum defects as a function of $1/[n-\delta^{(l,J)}]^{2}$, shown as Fig. 2(b). The difference between $nS_{1/2}$ and $nD_{J}$ is the result of a competition between core polarization and core penetration. Core polarization reduces the quantum defect, more so as the electron effective radius is reduced, i.e., low $n$. As $n$ increases, the core polarization effect decreases faster than the core penetration effect, which leads to a quantum defect dominated by core penetration.
Weber et al. [45] investigated the contribution from core polarization for the $nD_{J}$ series of Cs and concluded that it is too small to account for the non-monotonic behavior. We quantitatively studied the behavior by numerically calculating the contributions to the quantum defect’s leading term from core polarization and core penetration. Using the model potential in Ref. [51] in conjunction with the spin–orbit coupling effect, we calculated the expectation values of the energy due to core polarization and core penetration [24]. The numerically determined contributions to $\delta^{(l,J)}_{0}$ are presented in Table 2. For the $nD_{J}$ series, the contribution from core polarization is over 40%, much larger than the estimate in Ref. [45], possibly because the second-order dipole polarization corrections were not included [52]. If core polarization is neglected, the $nD_{J}$ quantum defects become larger as $n$ decreases, similar to $nS_{1/2}$. For the $nS_{1/2}$ series, the contribution from core polarization is less than 2.5%.
Table 3: Fine-structure interval parameters determined for $nD_{J}$ series both experimentally (Exp) and numerically (Num). The quoted uncertainties are the standard deviations of the fit.
$nD_{3/2}\rightarrow nD_{5/2}$ $\xi^{(l)}_{1}$ (kHz) $\xi^{(l)}_{2}$ (kHz) $\xi^{(l)}_{3}$ (kHz)
Exp [this work] 6.01624(22)×$\times 10^{10}$-3.38(24)×$\times 10^{10}$ 8(6)×$\times 10^{11}$
Exp [53] 6.02183(60)×$\times 10^{10}$-5.8(8)×$\times 10^{10}$ 00
Num [this work] 6.547×$\times 10^{10}$ 5.5 ×$\times 10^{10}$ 7.7×$\times 10^{11}$
We studied the fine-structure splitting of the $nD_{J}$ series. The fine-structure splitting, $\Delta_{\rm fs}$, can be expanded in terms of a power series in the effective principal quantum number, [$n-\tilde{\delta}^{(l,J=l\pm 1)}$], [53]:
$\Delta_{\rm fs}=\sum_{q=1}^{\infty}\frac{\xi^{(l)}_{q}}{[n-\tilde{\delta}^{(l,J=l\pm 1)}(n)]^{2q+1}}.$ (3)
The weighted average quantum defect, $\tilde{\delta}^{(l,J=l\pm 1)}(n)$, has the form
$\tilde{\delta}^{(l,J=l\pm 1)}(n)=w^{(l,J<l)}{\delta}^{(l,J<l)}(n)+w^{(l,J>l)}{\delta}^{(l,J>l)}(n).$ (4)
The weights for the fine-structure splitting, $w^{(l,J)}$, are equal to $1-\frac{|\langle{\textbf{L}}\cdot{\textbf{S}}\rangle^{(l,J)}|}{\hbar^{2}(l+1/2)}$, where $\langle{\textbf{L}}\cdot{\textbf{S}}\rangle^{(l,J)}=\frac{\hbar^{2}}{2}[J(J+1)-l(l+1)-\frac{3}{4}]$. For the $nD_{J}$ series, $w^{(2,3/2)}=0.4$ and $w^{(2,5/2)}=0.6$.
Figure 3: The fine-structure splitting for $nD_{J}$. The red dashed line shows the prediction using Eq. (3). The fitting residuals are given as the black circles in the inset. The fine-structure interval based on the fitting coefficients reported in Ref. [53] is plotted as the blue dashed line for comparison. The corresponding residuals in data in Ref. [53], are shown as grey squares in the inset.
Using the quantum defect values in Table 1 and Eq. (4), the measured fine-structure intervals were fit to Eq. (3) to extract $\xi^{(l)}_{q}$ up to $q=3$. The values of $\xi^{(l)}_{q}$ were reported in Ref. [53] based on a single fine-structure interval measurement. The parameters obtained from the measurements presented here are given in Table 3. The quality of the fit is evident in Fig. 3; the extracted parameters better describe the fine-structure splittings, $\Delta_{\rm fs}$, at low $n$. The fitting residuals are randomly distributed around zero as shown in the inset of Fig. 3.
The use of Eq. (3) to describe $\Delta_{\rm fs}$ had been proposed and used in previous work [53]. The physical meaning of $\xi^{(l)}_{q}$, have not been discussed to our knowledge. Here, we justify the expansion (3) by analytically deriving the spin–orbit coupling constant with hydrogenic wave functions, $\langle\frac{1}{r}\frac{{\rm d}V(r)}{{\rm d}r}{\textbf{L}}\cdot{\textbf{S}}\rangle$, for the atomic potential in the form of the sum of the Coulomb and polarization potentials, see Ref. [24]. Moreover, using the effective non-local model potential from Ref. [51] and the precise quantum defects, accurate wave functions and spin–orbit energies for the $nD_{5/2}$ and $nD_{3/2}$ series are determined. The numerical predictions $\Delta_{\rm fs}$ are obtained by taking the differences between the expectation values of the spin–orbit coupling operator of the two series [24]. They are fit to Eq. (3) to extract $\xi^{(l)}_{q}$, as shown in Table 3. The difference in the sign of $\xi^{(l)}_{2}$ may result from the model potential adopted for the numerical calculations. Importantly, the values of the leading coefficient for both the experimental and numerical approaches, $\xi^{(l)}_{1}$, are of similar order of magnitude. We note that the weighted average quantum defect, $\tilde{\delta}^{(l,J=l\pm 1)}(n)$, should be weighted by the total angular momenta of the $nD_{5/2}$ and $nD_{3/2}$ series, rather than being weighted equally as was done in Ref. [53].
Using our quantum defects and those of Ref. [49] for the $P$ states, we computed reduced electric-dipole matrix elements for $nP_{J}\leftrightarrow n^{\prime}D_{J^{\prime}}$ transitions of Cs. The calculated matrix elements for the low-lying Rydberg states agree within theoretical uncertainties, $<3\%$, with high-precision, relativistic, all-order many-body calculations of Safronova et al. [21], which are computationally sophisticated and resource intensive to perform. With accurate quantum defects, garnered from high $n$ states, accurate wave functions and matrix elements can be obtained for all $n$, notably for low $n$ where the spectroscopy is difficult to carry-out because of the energy separations between different states. The full comparison table of the matrix elements for the $nP_{J}\leftrightarrow n^{\prime}D_{J^{\prime}}$ transitions is presented in the supplemental material, Table S3 [24].
In summary, we presented high precision, absolute frequency measurements of Cs energy levels from the $|6S_{1/2},F=3\rangle$ hyperfine ground state to $|nS_{1/2}$ $(n=23-90)\rangle$, $|nD_{3/2}$ $(n=21-90)\rangle$, and $|nD_{5/2}$ $(n=21-90)\rangle$ Rydberg states with an accuracy of $<72\,\rm kHz$. We investigated the validity of the modified Ritz expansion and conclude that a modified Ritz expansion with only even quantum defect terms can accurately describe the energy levels of the Cs $nS_{1/2}$ and $nD_{J}$ states. By fitting the absolute-frequency measurements to the modified Ritz formula, Eq. (1), we determine the quantum defects of the Cs $nS_{1/2}$, $nD_{3/2}$, and $nD_{5/2}$ series and the ionization energy. The ionization energy was found to be $31406.46775148(14)~{}\rm{cm}^{-1}$. At the current level of precision, we find the quantum defect expansion can be terminated at the third order. We separated the contributions from core polarization and core penetration. For the $nD_{J}$ series, the contribution from core polarization is about 40%, while for the $nS_{1/2}$ series, the contribution is less than 2.5%. Using improved Cs Rydberg wave functions derived from the quantum defects, we recalculated reduced electric-dipole matrix elements for the $nP_{J}\leftrightarrow n^{\prime}D_{J^{\prime}}$ transitions. The obtained matrix elements are found to be in agreement with high level, many-body relativistic calculations at low $n$, where our measurements were extrapolated.
Some aspects of prior work that were not accurate have been corrected, such as the form of the fine-structure formula and the number of quantum defect parameters required to describe the spectrum at our level of precision. The accuracy of the wave functions derived from the energy levels of the Cs atom promise to enable a host of important calculations that are not possible to do accurately now, such as those in few-body physics involving Rydberg states and their interatomic potentials. Furthermore, the accuracy of the dipole moments between different Rydberg states pushes Rydberg-atom-based sensors towards becoming a primary standard for radio frequency fields. The current state of art transition dipole moments for Cs Rydberg RF transitions are accurate to within 0.5 % [54]. Precise knowledge of the atomic wave functions is necessary not only to confirm this assertion but to improve the accuracy. Future work will consist of constructing a new model potential [51] for Cs and the verification of the transition dipole moments.
Acknowledgements.
This work has been supported by The National Research Council Internet of Things: Quantum Sensors Challenge program through Contract No. QSP-058-1. The calculations were supported at ITAMP by a grant from the US National Science Foundation.